hallm:

The area of a rectangle is 72 cm². The length of the rectangle is 6 cm longer than the width. What is the width of the rectangle? 2 cm 6 cm 8 cm 10 cm

Shadow:

Hello @hallm

hallm:

hey @Shadow

Shadow:

Do you have an idea where you might start in order to solve this problem?

hallm:

no sir im reallt confused

Shadow:

Do you know the formula for the area of a rectangle?

hallm:

umm length time width right?

Shadow:

Yes

Shadow:

Write out the formula for me

hallm:

wouldnt it be 72 times 72 because of the squared sign and then divide it by 6 maybe....?

hallm:

wouldnt it be you times 72 by 72 because of the squared then divide it by 6 maybe...?

Shadow:

The cm^2 just means that they are square centimeters since we’re dealing with a 2 D object. So it is 72 two by two centimeter squares.

Shadow:

It will make more sense when we go through the math. The formula for the area of a rectangle is: A = length times width A = l * w

Shadow:

They have given us two pieces of information. The area is 72cm^2 and the length is 6 cm longer than the width. The first bit of information is easy to account for. 72 = l * w But what about the second? Well if the length is 6 cm longer than the width then we should be able to represent it was l= w + 6

Shadow:

This is crucial as this means we can substitute w + 6 in for l, meaning we only have one variable. Thus we can solve for w. 72 = (w + 6) * w Do you know what to do from here?

hallm:

then you divide both 6 by 6and 72 by 6?

Shadow:

Well things will get a bit messy if we do it that way. What’s more easier?

hallm:

hmm well maybe umm i dont know im still confused

hallm:

i feel kinda dumb for not knowing this

Shadow:

Well we all didn’t know how to do this until we found out, so don’t feel bad

hallm:

yes sir maybe we can do the pemdas method and umm then divide?

Shadow:

Try the distributive property.

hallm:

ok

hallm:

i got 6

Shadow:

Correct

Shadow:

You can check your work by submitting the value you got for w back into l = w + 6 then seeing if the length and width values come out to 72.

hallm:

thank you so much you have really helped me

Shadow:

No problem, glad I could help.

hallm:

what grade are you in ?

Shadow:

Freshman in college

hallm:

wooowww thats so cool

hallm:

what do you study

Shadow:

My majors are psychology and political science.

hallm:

wow that seems really hard im just a freshman in highschool

Shadow:

They are social sciences so they are a bit on the easier side on the spectrum of all possible majors but they do get increasingly more difficult. I found out what I wanted to major in through classes and teachers I had in high school as well as extracurricular. So I would highly encourage you to explore your interests, talk to teachers if you find something interesting, and always check out what sports/clubs are going on. You may never know what you’ll find that you may enjoy.

hallm:

well school isnt going as well as i would have planed it is honestly really hard

Shadow:

Well whenever it does get difficult, just be sure to always ask for help. Being honest that you can’t accomplish something doesn’t mean that you are dumb, it means that you are wise. By choosing to seek out help you can learn how to do a concept and this become more knowledgeable.

hallm:

thank you